OpenStudy (anonymous):

Use properties of trigonometric functions to find the exact value of the expression: (sin (-31)/cos (391)) + tan (211)

OpenStudy (blockcolder):

\[\cos391=\cos31=\cos(-31)\\ \tan211=\tan31\\ \frac{\sin-31}{\cos-31}+\tan31=\tan-31+\tan31=-\tan31+\tan31=0\]

OpenStudy (anonymous):

thank you!